OpenStudy (anonymous):

Trying to figure out the method for solving [(x-9)(x+1)]/(x-2) greater than or equal to 0

OpenStudy (saifoo.khan):

\[\frac{(x-9)(x+1)}{x-2} \ge 0 \]

OpenStudy (saifoo.khan):

\[(x-9)(x+1)\ge 0 \]

OpenStudy (saifoo.khan):

\[x \ge 9 , x \le - 1\]

OpenStudy (anonymous):

Thanks. How do you express that in interval notation?

myininaya (myininaya):

saifoo you can't do that since x-2 is also sometimes negative

OpenStudy (saifoo.khan):

myininya, quick question, is -3 and -6 complex?

myininaya (myininaya):

\[f(x)=\frac{(x-9)(x+1)}{x-2} \] f=0 when x=9, =-1 f is undefined at x=2 so we need to look around each of these numbers ------|--------|----------|--------- -1 2 9 Test all 4 of these intervals.

myininaya (myininaya):

yes anything in a+bi form is complex b can be 0

myininaya (myininaya):

where a and b are real

OpenStudy (saifoo.khan):

O I C.

OpenStudy (saifoo.khan):

So my answer is incorrect right?

myininaya (myininaya):

\[f(-2)=\frac{(-2-9)(-2+1)}{-2-2}=\frac{(-)(-)}{-}=-<0\] \[f(0)=\frac{(-)(+)}{(-)}=(+)>0\] \[f(3)=\frac{(-)(+)}{(+)}=(-)<0\] \[f(10)=(+)>0\] so f>=0 when...

myininaya (myininaya):

\[x \in [-1,2) \cup [9,\infty)\]